OpenStudy (anonymous):

a quadratic function is defined by y=3x^2+4x-2. a linear function is defined by y=mx-5. what value(S) of the slope of the line would make it tangent to the parabola

OpenStudy (anonymous):

y=3x²+4x-2 y' is the slope of any tangent line y'=6x+4 the point of tangency must satisfy both eqaution of the parabola and the the equation of the tangent line. y=mx-5 3x²+4x-2 = (6x+4)x -5 solve for x 3x²+4x-2 = 6x²+4x - 5 -3x² = -3 x² = 1 x = ±1 for x=1, y=3(1)²+4(1)-2 y=3+4-2 y=5 subs to y=mx-5 5=m(1)-5 m=10 for x=-1, y=3(-1)²+4(-1)-2 y=3-4-2 y=-3 subs to y=mx-5 -3=m(-1)-5 m=-2 answers: m = 10, -2

OpenStudy (anonymous):

how did you get 6x+4?

OpenStudy (anonymous):

hold on what math are you in?

OpenStudy (anonymous):

grade 11 functions

OpenStudy (anonymous):

ok good. making sure i didnt do it wrong. My teacher went over this exact problem today and i wrote it down. I didnt have time to ask how he actually got the 6x+4 cause of the bell but i am going to tomorrow

OpenStudy (anonymous):

oh ok